OpenStudy (anonymous):

Will medal and fan please help!!! Noelle needs to have her computer fixed. One company charges a 25$ estimation fee plus 45$ per hour. Another company charges a 10$ estimation fee plus $60 per hour, Find the number of hours for which the two costs would be the same.

OpenStudy (anonymous):

We have two known numbers without any further need of changing (estimation fee) and the time fee. So, the costs are, as formula: for company A \[25+45x =y\] for company B \[10+60x=y\] x is the hours and y is the final cost when they are equal. Does that help you at all? @Twix47

OpenStudy (anonymous):

yeah thanks

OpenStudy (anonymous):

So, the easiest is to set them as equal to each other: \[25+45x=10+60x\] Now simplify and solve for x.